OpenStudy (anonymous):

just want to know if I did this problem right. In triangle ABC, a = 9, c = 5, and B = 120°. Find b2 I got 128.5

OpenStudy (anonymous):

@jagr2713

OpenStudy (campbell_st):

well if you are using the cosine rule its \[b^2 = 9^2 + 5^2 - 2 \times 5 \times 9 \times \cos(120)\]

OpenStudy (campbell_st):

so I think your answer is incorrect... \[2 \times 5 \times 9 \times \cos(120) = -45\] so the problem is really \[b^2 = 9^2 + 5^2 - (-45)\]

OpenStudy (anonymous):

so 151 @campbell_st ??

OpenStudy (campbell_st):

that's what I got

OpenStudy (anonymous):

okay, can you help me with another one please? @campbell_st

OpenStudy (campbell_st):

sorry, I'm about to go to class.